OpenStudy (anonymous):

Write the equation of the given line in slope-intercept form. Coordinates are (-1, 2) & (1, -4) A. y= 3x - 1 B. y = -3x - 1 <---- C. y = -3x + 1 D. y = -1/3 x - 1 THE ANSWER IS B

pooja195 (pooja195):

Hi welcome to openstudy :) Start by finding the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{-4-2}{1--1 }=slope\] \[\LARGE \frac{-4-2}{1+1 }=slope\] Find the slope

OpenStudy (anonymous):

y=mx+b m= slope

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

slope is y over x (y/x)

OpenStudy (anonymous):

is it -6/2?

OpenStudy (mathmale):

pooja's formula for the slope is the more relevant one here.

OpenStudy (mathmale):

Vicky: I'd be happy to give you feedback if yo u show your work.

pooja195 (pooja195):

Yes what is -6 divided by 2? we can always simplify the slope

pooja195 (pooja195):

^If it's possible to

OpenStudy (anonymous):

well i just finished your equation...i got 3/ 1?

OpenStudy (anonymous):

-

OpenStudy (mathmale):

You've reduced the fraction representing the slope, but you have not "finished your equation."

pooja195 (pooja195):

When we divide by ONE negative our solution will include a - so 3 is correct but it would be -3

OpenStudy (anonymous):

oooh ok i see

OpenStudy (mathmale):

the desired equation is y=mx+b. Supposedly you've found the slope, m.

pooja195 (pooja195):

^use that y=-3x+b

OpenStudy (mathmale):

Here pooja has substituted -3 for the slope, m.

OpenStudy (mathmale):

How would you now find the value of the y-intercept, b? Hint: you started out with two points.

OpenStudy (anonymous):

yeah, it's not c

OpenStudy (anonymous):

basically...you now have the equation y=-3x+b

OpenStudy (anonymous):

b = y-intercept or when x=0

OpenStudy (anonymous):

A simple way of solving a problem like this, if you were given multi choices is to just plug in the the x and y cords into the equation and see if it works out

OpenStudy (anonymous):

Im still lost

OpenStudy (anonymous):

Wait nevermind i think i got it lol sorry

OpenStudy (anonymous):

Do you want to know the simple way of solving the problem or do you want to know how to construct the equation from the 2 sets of points

OpenStudy (anonymous):

Its A right?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

b sorry

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Sorry it took me so long to figure it out! I really appreciate your patience :)